A cable company charges $75 for installation plus $20 per month. Another cable company offers free installation but charges $35 per month. For how many months of cable service would the total cost from either company be the same?
a)2 b) 3 c) 5 d) 7

Answer by stanbon(48568)

A cable company charges $75 for installation plus $20 per month. Another cable company offers free installation but charges $35 per month. For how many months of cable service would the total cost from either company be the same?
a)2 b) 3 c) 5 d) 7
----------------------------
1st Co Eq: A(x) = 75 + 20x where x is # of months of service
-------
2nd Co Eq; B(x) = 35x where x is # of months of service
---------------------------------
Equation:
When does A(x) = B(x)
75+20x = 35x
15x = 75
x = 5 months
==================
Cheers,
Stan H.